# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Introduction
where $\bm{\alpha}_{0_{K\times 1}},\bm{\alpha}_{1_{K\times 1}},$ and $\bm{\gamma}_{L\times 1}$ are the dual variables associated with the balancing constraints in Equations (5b)-(5c), $\rho(t)=t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)$ is strictly concave, and h(t)=D(1t)h(t)=D(1-t). Furthermore, the primal solution $w_{i}^{*}=Z_{i}w^{*}_{1i}+(1-Z_{i})w^{*}_{0i}$ satisfies $w^{*}_{1i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{1}^{T}c(\bm{X}_{i})+\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}\>\>\text{and}\>\>w^{*}_{0i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X}_{i})-\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}$ where $\hat{\bm{\alpha}}_{0},\hat{\bm{\alpha}}_{1},\hat{\bm{\gamma}}$ is the solution to the dual problem.
The proof of this theorem is in supplementary Section 1. In the following sections, we discuss the implications of Problem (5) in terms of both positivity and the interpretation of the resulting estimator while drawing comparisons to direct balancing weights, minimal weights, and overlap weights.
@@
The construction of Problem (5) is primarily motivated by the challenges posed by lack of overlap on causal effect estimation, generally, and direct balancing weights, specifically. In the following section, we show that Problem (5) corresponds to a relaxed positivity assumption where there is a solution to Problem (5) even when the standard positivity assumption does not hold. Furthermore, we discuss the restrictions placed on the implied propensity score model of Problem (5) and how this restriction connects to the relaxed positivity assumption. We additionally connect these results to the positivity assumptions and implied propensity score models of direct balancing, minimal, and overlap weights.
While positivity is required for the existence of a solution to Problem (3) (Zhao and Percival,, 2017), by not balancing $g(\bm{x})$ to the target population, the weights derived through Problem (5) exist under a relaxed positivity assumption:
@@
The following proposition states that this assumption is sufficient for the existence of the weights in Problem (5) as $n\to\infty$.
@@
We leave the proof of this proposition to supplementary Section 1, however here we provide a brief intuitive discussion of how Problem (5) often yields a solution in finite samples when Problem (3) does not. The weights in Problem (3) exist if the convex hulls generated by $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=1}$ and $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=0}$ contain $\frac{1}{n}\sum_{i=1}^{n}\{c(\bm{X}_{i}),g(\bm{X}_{i})\}$ (Zhao and Percival,, 2017). In contrast, the weights in Problem (5) exist if 1) the convex hulls of $\{c(\bm{X}_{i})\}_{Z_{i}=1}$ and $\{c(\bm{X}_{i})\}_{Z_{i}=0}$ contain $\frac{1}{n}\sum_{i=1}^{n}c(\bm{X}_{i})$; and 2) $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=1}\cap\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=0}\neq\emptyset$. Thus, as long as any $J$ dimensional vector is in the intersection of the convex hulls, rather than the $J$ dimensional sample mean, the existence of a solution is guaranteed. This is a substantial relaxation of the requirements for weight existence, especially for scenarios where one treatment group is much smaller than the other. Returning to Figure 1, the sample mean of $\{c(\bm{x}),g(\bm{x})\}$ is not in the intersection of the two convex hulls and thus there is no solution for direct balancing weights as previously discussed. However, there are many observations in the intersection of the two convex hulls. Further, when examining $c(\bm{x})$, its sample mean is within the observed range of $c(\bm{x})$ values for both treatment groups. Thus, there is a solution to our proposed Problem (5), and in fact the corresponding weighted sample mean given by $\{w_{i}^{*}\}_{i=1}^{n}$ is within the intersection of the treated and control convex hulls.
The relaxed positivity assumption for Problem 5 is directly related to the propensity score model implied by Problem 5. Taking the expectation of population version of the loss function in Equation (6) with respect to $\bm{x}$ yields the following first order conditions,
@@
Thus, the weights $w_{0}^{*}$ and $w_{1}^{*}$ correspond to IPW for a specific propensity score model, which we reference as the implied propensity score model of Problem (5). In this implied propensity score model, the dual variables $\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}$ can be considered as the model coefficients. Furthermore, $\rho^{\prime}(\cdot)$ may be regarded as the link function of the generalized linear propensity score model, where the selected measure of dispersion (i.e., $D(\cdot)$) determines this link function (Wang and Zubizarreta,, 2020).
Problem (5) restricts the coefficients corresponding to $g(\bm{x})$ in the models for $\{e(\bm{x})\}^{-1}$ and $\{1-e(\bm{x})\}^{-1}$ to be opposites of each other. This ensures that $w^{*}_{0}=\{1-w_{1}^{*^{-1}}\}^{-1}$ such that only a single propensity score model is specified by Problem (5). However, there there are substantial restrictions on the propensity score model that satisfies conditions (7a)-(7c). Specifically, $\text{Var}(Z|\bm{X}=\bm{x})$ is restricted to be a function of only $c(\bm{x})$ for specific dispersion measures (see supplementary Section 2 for further discussion). Since this generally does not hold when $E[Z|\bm{X}=\bm{x}]$ is a function of both $c(\bm{x})$ and $g(\bm{x})$, this model will only hold when $g(\bm{x})$ takes a specific functional form. However, it is explicitly through this restriction on the propensity score model that there is a solution to Problem (5) in scenarios with a lack of overlap, as overlap is improved through a reduction in the ability to predict treatment assignment (Clivio et al.,, 2024). Rather than attempt to specify a correct implied propensity score model, Problem (5) imposes restrictions on the propensity score model to improve overlap and yield a solution to the optimization problem.
@@
Problem (5) corresponds to a single propensity score model, while Problem (3) implies two incompatible propensity score models due to the separability of the optimization problem by treatment group. By implying a single propensity score model, rather than two incompatible propensity score models, the weights derived by our proposed method maintain the standard philosophical interpretation of inverse probability weights that direct balancing weights lack for the ATE. Furthermore, this allows for the doubly robust property (i.e., the estimator is consistent if either the implied propensity score model or outcome model is properly specified) and semiparametric efficiency of our method given the correct model specification that direct balancing weights for the ATE do not satisfy (see Section 4 for further discussion of these properties). Minimal weights also correspond to two incompatible propensity score models but place an l1 penalty on the dual variables (i.e., propensity score model coefficients). While the restrictions imposed on the implied propensity score model of Problem (5) are comparably less intuitive, by restricting $\text{Var}(Z|\bm{X}=\bm{x})$ rather than $E[Z|\bm{X}=\bm{x}]$ our derived weights are able to achieve exact finite sample balance, unlike minimal weights. Furthermore, reducing the variance of the derived weights is key to reducing estimator variance, which restricting $\text{Var}(Z|\bm{X}=\bm{x})$ yields; thus, our proposed method precisely restricts the aspect the propensity score model that primarily influences estimator variance.
When $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ in Problem (5), Assumption Assumption 2: Relaxed positivity becomes: “There exists an $\bm{x}^{*}$ such that $0<\text{Pr}\{Z_{i}|c(\bm{X}_{i})=c(\bm{x}^{*}),g(\bm{X}_{i})=g(\bm{x}^{*})\}<1$” which is comparable to the lack of a positivity assumption for overlap weights. Thus, Assumption Assumption 2: Relaxed positivity is an intermediate between Assumption $2$ and no positivity assumption; as more covariate functions are added to the $g(\bm{x})$ set, Assumption Assumption 2: Relaxed positivity becomes more relaxed. To further connect overlap weights and Problem (5), when $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ the first order conditions in (7) simplify to $e(\bm{x})/\{1-e(\bm{x})\}=w_{0}^{*}/w_{1}^{*}$. While the exact solution to the dual problem in this case will depend on the dispersion measure, overlap weights satisfy this condition. Therefore, Problem (5) with $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ can be considered the direct balancing equivalent of overlap weights; however, Problem (5) will always yield exact balance while overlap weights only yields exact balance when the propensity score is estimated with a logistic regression.
@@
While Problem (5) yields a solution when there is a lack of overlap, in doing so the population of $g(\bm{x})$ is modified away from the ATE target population, potentially resulting in estimator bias with respect to the ATE (termed bias due to estimand mismatch in Barnard et al., (2025)). In the following sections, we provide model- and design- based perspectives on the conditions under which our proposed estimator has minimal error with respect to the ATE. In addition, we provide a characterization of the modified population and estimand that $\tau_{w^{*}}$ targets when $\tau_{w^{*}}$ does not have minimal error with respect to the ATE. To discuss when our proposed estimator has minimal error with respect to the ATE, we first explore the following error decomposition for a normalized weighted estimator for the ATE,
$\hskip-14.45377pt\hat{\tau}_{w}-\tau$ $=\frac{1}{n}\sum_{i=1}^{n}\{w(\bm{X}_{i})Z_{i}\mu_{Z_{i}}(\bm{X}_{i})-\mu_{1}(\bm{X}_{i})\}-\frac{1}{n}\sum_{i=1}^{n}\{w(\bm{X}_{i})(1-Z_{i})\mu_{Z_{i}}(\bm{X}_{i})-\mu_{0}(\bm{X}_{i})\}$ (8a)
@@
$+\frac{1}{n_{1}}\sum_{i=1}^{n}w_{1}(\bm{X}_{i})\epsilon_{i}Z_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}w_{0}(\bm{X}_{i})\epsilon_{i}(1-Z_{i}),$ (8c)
@@
To better understand the conditions under which our balancing equations are justified, consider the case where $\mu_{z}(\bm{x})$ is a linear function of $c(\bm{x})$ and $g(\bm{x})$ such that $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. When weights are derived through Problem (3), term (8a) will be zero for the corresponding estimator. However when weights are derived with our proposed Problem (5), term (8a) simplifies to $\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{1}^{T}g(\bm{X}_{i})-\bm{\lambda}_{0}^{T}g(\bm{X}_{i})$ implying that term (8a) is zero when $\bm{\lambda}_{1}=\bm{\lambda}_{0}$. Thus, term (8a) for our proposed estimator is zero when $\mu_{z}(\bm{x})=\bm{\beta}_{z}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ and $\tau(\bm{x})=(\bm{\beta}_{1}^{T}-\bm{\beta}_{0}^{T})c(\bm{x})$. These models imply that our proposed estimator has minimal error with respect to the ATE when the set $c(\bm{x})$ contains all treatment effect modifiers. Thus, as the set $c(\bm{x})$ gets smaller and correspondingly the set $g(\bm{x})$ gets larger, our proposed estimator requires a stronger assumption about the number of effect modifiers to be justified for the ATE. However, this is a relaxation of the assumption $\tau(\bm{x})=\tau$ required for the overlap weights estimator to be unbiased for the ATE.

### v2

@@ Introduction
where $\bm{\alpha}_{0_{K\times 1}},\bm{\alpha}_{1_{K\times 1}},$ and $\bm{\gamma}_{L\times 1}$ are the dual variables associated with the balancing constraints in Equations (5b)-(5c), $\rho(t)=t-t(h^{\prime})^{-1}(t)+h[(h^{\prime})^{-1}(t)]-h(1)$ is strictly concave, and h(t)=D(1t)h(t)=D(1-t). Furthermore, the primal solution $w_{i}^{*}=Z_{i}w^{*}_{1i}+(1-Z_{i})w^{*}_{0i}$ satisfies $w^{*}_{1i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{1}^{T}c(\bm{X}_{i})+\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}\>\>\text{and}\>\>w^{*}_{0i}=\rho^{\prime}\{\hat{\bm{\alpha}}_{0}^{T}c(\bm{X}_{i})-\hat{\bm{\gamma}}^{T}g(\bm{X}_{i})\}$ where $\hat{\bm{\alpha}}_{0},\hat{\bm{\alpha}}_{1},\hat{\bm{\gamma}}$ are the solution to the dual problem.
The proof of this theorem is in supplementary Section 1. In the following sections, we discuss the implications of Problem (5) in terms of positivity, the interpretation of the resulting estimator, and related estimation problems while drawing comparisons to direct balancing weights, minimal weights, and overlap weights.
@@
Problem (5) is primarily motivated by the challenges posed by lack of overlap on causal effect estimation, generally, and direct balancing weights, specifically. In the following section, we show that a solution to Problem (5), with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$, exists under a relaxed positivity assumption. Furthermore, we determine the propensity score model implied by Problem (5) and describe how this model connects to the relaxed positivity assumption and other methods. While positivity is required for the existence of a solution to Problem (3) with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ (Zhao and Percival, 2017), by not balancing $g(\bm{x})$ to the target population, the weights derived through Problem (5) exist under a relaxed positivity assumption:
@@
The following proposition states that this assumption is sufficient for the existence of the weights in Problem (5) with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ as $n\to\infty$.
@@
We leave the proof of this proposition to supplementary Section 1. When considering a finite sample, the weights in Problem (3) exist if the convex hulls generated by $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=1}$ and $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=0}$ contain $\frac{1}{n}\sum_{i=1}^{n}\{c(\bm{X}_{i}),g(\bm{X}_{i})\}$ (Zhao and Percival, 2017). In contrast, the weights in Problem (5) exist if 1) the convex hulls of $\{c(\bm{X}_{i})\}_{Z_{i}=1}$ and $\{c(\bm{X}_{i})\}_{Z_{i}=0}$ contain $\frac{1}{n}\sum_{i=1}^{n}c(\bm{X}_{i})$; and 2) $\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=1}\cap\{c(\bm{X}_{i}),g(\bm{X}_{i})\}_{Z_{i}=0}\neq\emptyset$. This is a substantial relaxation of the requirements for weight existence, especially for scenarios where one treatment group is much smaller than the other. Returning to Figure 1, the sample mean of $\{c(\bm{x}),g(\bm{x})\}$ is not in the intersection of the two convex hulls and thus there is no solution to Problem (3) as previously discussed. However, there are many observations in the intersection of the two convex hulls. Furthermore, when examining $c(\bm{x})$, its sample mean is within the observed range for both treatment groups. Thus, there is a solution to our proposed Problem (5), and in fact the corresponding weighted sample mean given by $\{w_{i}^{*}\}_{i=1}^{n}$ is within the intersection of the treated and control convex hulls.
The relaxed positivity assumption for Problem 5 is directly related to the propensity score model implied by Problem 5. Taking the expectation of population version of Equation (6) with respect to $\bm{x}$ yields the following first order conditions,
@@
Thus, the weights $w_{0}^{*}$ and $w_{1}^{*}$ correspond to IPW for a specific propensity score model, which we reference as the implied propensity score model of Problem (5). In this implied propensity score model, the dual variables $\bm{\alpha}_{0},\bm{\alpha}_{1},\bm{\gamma}$ can be considered as the model coefficients. Furthermore, $\rho^{\prime}(\cdot)$ may be regarded as the link function of the generalized linear propensity score model, where $D(\cdot)$ determines this link function (Wang and Zubizarreta, 2020).
Problem (5) restricts the coefficients corresponding to $g(\bm{x})$ in the models for $\{e(\bm{x})\}^{-1}$ and $\{1-e(\bm{x})\}^{-1}$ to be opposites of each other. This ensures that $w^{*}_{0}=\{1-w_{1}^{*^{-1}}\}^{-1}$ such that only a single propensity score model is specified by Problem (5). However, there are substantial restrictions on the propensity score model that satisfies conditions (7a)-(7c). Specifically, $\text{Var}(Z|\bm{X}=\bm{x})$ is restricted to be a function of only $c(\bm{x})$ for specific dispersion measures (see supplementary Section 2 for further discussion). Since this generally does not hold when $E[Z|\bm{X}=\bm{x}]$ is a function of both $c(\bm{x})$ and $g(\bm{x})$, this model will only hold when $g(\bm{x})$ takes a specific functional form. However, it is explicitly through this restriction on the propensity score model that there is a solution to Problem (5) in scenarios with a lack of overlap, as overlap is improved through a reduction in the ability to predict treatment assignment (Clivio et al., 2024). Rather than attempt to specify a correct implied propensity score model, Problem (5) imposes restrictions on the propensity score model to improve overlap and yield a solution to the optimization problem.
@@
Problem (5) corresponds to a single propensity score model, while Problem (3) implies two incompatible propensity score models due to the separability of the optimization problem by treatment group. By implying a single propensity score model, the weights derived by our proposed method maintain the standard philosophical interpretation of IPW that direct balancing weights lack for the ATE. Furthermore, this allows for the doubly robust property (i.e., the estimator is consistent if either the implied propensity score model or outcome model is properly specified) and semiparametric efficiency of our method given the correct model specification (see Section 4 for further discussion of these properties). Minimal weights also correspond to two incompatible propensity score models but additionally place an L1 penalty on the dual variables (i.e., propensity score model coefficients). While the restrictions imposed on the implied propensity score model of Problem (5) are comparably less intuitive, by restricting only $\text{Var}(Z|\bm{X}=\bm{x})$ rather than both $E[Z|\bm{X}=\bm{x}]$ and $\text{Var}(Z|\bm{X}=\bm{x})$ our derived weights are able to achieve exact finite sample balance. Furthermore, in restricting $\text{Var}(Z|\bm{X}=\bm{x})$ our proposed method precisely restricts the aspect the propensity score model that primarily influences estimator variance.
When $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ in Problem (5), Assumption Assumption 2: Relaxed positivity becomes: “There exists an $\bm{x}^{*}$ such that $0<\text{Pr}\{Z_{i}|c(\bm{X}_{i})=c(\bm{x}^{*}),g(\bm{X}_{i})=g(\bm{x}^{*})\}<1$” which is comparable to the lack of a positivity assumption for overlap weights. Thus, Assumption Assumption 2: Relaxed positivity is an intermediate between Assumption $2$ and no positivity assumption; as more covariate functions are added to the $g(\bm{x})$ set, Assumption Assumption 2: Relaxed positivity becomes more relaxed. To further connect overlap weights and Problem (5), when $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ the first order conditions in (7) simplify to $e(\bm{x})/\{1-e(\bm{x})\}=w_{0}^{*}/w_{1}^{*}$. While the exact solution to the dual problem will depend on the specific dispersion measure, overlap weights satisfy this condition. Therefore, Problem (5) with $c(\bm{x})=\emptyset$ and $g(\bm{x})=b(\bm{x})$ can be considered the direct balancing equivalent of overlap weights; however, Problem (5) will always yield exact balance while overlap weights only yields exact balance when the propensity score is estimated with logistic regression.
@@
While Problem (5) yields a solution when there is a lack of overlap, in doing so the population of $g(\bm{x})$ is modified away from the ATE target population, potentially resulting in estimator bias with respect to the ATE (termed bias due to estimand mismatch in Barnard et al. (2025)). Thus, we provide model- and design- based perspectives on the conditions under which our proposed estimator has minimal error with respect to the ATE. In addition, we provide a characterization of the modified population and estimand that $\hat{\tau}_{w^{*}}$ targets when $\hat{\tau}_{w^{*}}$ does not have minimal error with respect to the ATE. We first note that for any weighted estimator for the ATE,
$\hskip-14.45377pt\hat{\tau}_{w}-\tau$ $=\frac{1}{n}\sum_{i=1}^{n}\{w_{i}Z_{i}\mu_{Z_{i}}(\bm{X}_{i})-\mu_{1}(\bm{X}_{i})\}-\frac{1}{n}\sum_{i=1}^{n}\{w_{i}(1-Z_{i})\mu_{Z_{i}}(\bm{X}_{i})-\mu_{0}(\bm{X}_{i})\}$ (8a)
@@
$+\frac{1}{n}\sum_{i=1}^{n}w_{i}\epsilon_{i}Z_{i}-\frac{1}{n}\sum_{i=1}^{n}w_{i}\epsilon_{i}(1-Z_{i}),$ (8c)
@@
To better understand the conditions under which our balancing equations are justified, consider the case where $\mu_{z}(\bm{x})$ is a linear function of $c(\bm{x})$ and $g(\bm{x})$ such that $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. When weights are derived through Problem (3), term (8a) will be zero for the corresponding estimator. However when weights are derived with Problem (5), term (8a) simplifies to $\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{1}^{T}g(\bm{X}_{i})-\bm{\lambda}_{0}^{T}g(\bm{X}_{i})$ such that term (8a) is zero when $\bm{\lambda}_{1}=\bm{\lambda}_{0}$. Thus, term (8a) is zero when $\mu_{z}(\bm{x})=\bm{\beta}_{z}^{T}c(\bm{x})+\bm{\lambda}^{T}g(\bm{x})$ and $\tau(\bm{x})=(\bm{\beta}_{1}^{T}-\bm{\beta}_{0}^{T})c(\bm{x})$. These models imply that our proposed estimator has minimal error with respect to the ATE when the $c(\bm{x})$ set contains all treatment effect modifiers. Thus, as the set $c(\bm{x})$ gets smaller and correspondingly the set $g(\bm{x})$ gets larger, our proposed estimator requires a stronger assumption about the number of effect modifiers to be justified for the ATE. However, this is a relaxation of the assumption $\tau(\bm{x})=\tau$ required for the overlap weights estimator to be consistent for the ATE.